# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ B.2 Proofs of Section 3’s Main Results
Combining these observations, we conclude that no creator $i$ has a profitable deviation from $(x_{i}^{\star},1)$, making $(\mathbf{x}^{\star},\mathbf{1})$ a PNE. Moreover, uniqueness follows from the observation that any FSE must be an ESE when restricted to the subgame with $\mathbf{s}=\mathbf{1}$. Since Proposition 3 establishes that the ESE is unique, there can be no other FSE. This completes the proof of Theorem 5. ∎
@@ Appendix C Omitted Proofs From Section 4
This section is divided into two parts. The first part is mainly technical, and its purpose is to provide the tools needed for the proof of Theorem 6, which will be the focus of the second part. As was discussed in the proof sketch of Theorem 6, our approach to analyzing the algorithm’s approximation guarantees includes three main steps:
@@
    First, we demonstrate that the ESE quality function $\mathcal{X}(\cdot)$ is Lipschitz continuous and bound its Lipschitz constant.
@@
    Second, we prove that the marginal gain a player can get by deviating from a full sharing profile is Lipschitz continuous as well and bound its constant.
@@
    Lastly, we show that the platform’s revenue is Lipschitz continuous as well, and bound its constant.
@@
$x_{i}^{\min}\coloneqq\left(\frac{dc_{i}}{dx_{i}}\right)^{-1}\left(\mu\gamma\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\right).$
C.1 Auxiliary Lemmas and Propositions for Theorem 6’s Proof
@@
Fix arbitrary $\rho$. Then, for every creator $i\in[n]$, it holds that $\mathcal{X}(\rho)_{i}\geq x_{i}^{\min}$.
Proof of Lemma 14.
The first-order condition (FOC) for creator $i$ in the ESE, given by the derivative of its utility function with respect to $x_{i}$, is:
$\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}\left[(\gamma-1)\mathcal{X}(\rho)_{i}+\left\|\mathcal{X}(\rho)\right\|\right]=c_{i}^{\prime}(\mathcal{X}(\rho)_{i}).$
Fix any creator $i$ and rearrange the LHS of its FOC:
$c_{i}^{\prime}(\mathcal{X}(\rho)_{i})$ $=\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}\left[(\gamma-1)\mathcal{X}(\rho)_{i}+\left\|\mathcal{X}(\rho)\right\|\right]$
$=\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\left[1+(\gamma-1)\frac{\mathcal{X}(\rho)_{i}}{\left\|\mathcal{X}(\rho)\right\|}\right].$
Notice that the function $f(y)=1+(\gamma-1)y$ is decreasing in $y$ for $\gamma<1$. Therefore, to lower bound $c_{i}^{\prime}(\mathcal{X}(\rho)_{i})$, we can evaluate $\left[1+(\gamma-1)\frac{\mathcal{X}(\rho)_{i}}{\left\|\mathcal{X}(\rho)\right\|}\right]$ at the maximum value of $\frac{\mathcal{X}(\rho)_{i}}{\left\|\mathcal{X}(\rho)\right\|}$, which is $1$. Using this observation, we get:
$c_{i}^{\prime}(\mathcal{X}(\rho)_{i})\geq\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\gamma.$
To further lower bound this expression, we can use the fact that $\left\|\mathcal{X}(\rho)\right\|\leq\sum_{j=1}^{n}x_{j}^{\max}$ for every $\rho\in[0,1]$ (since $\mathcal{X}(\rho)_{i}\leq x_{i}^{\max}$ for all $i$). Then, using the fact that $\gamma\leq 1$, we have:
$c_{i}^{\prime}(\mathcal{X}(\rho)_{i})\geq\frac{\mu(1+\alpha\rho)}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma\geq\mu\gamma\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}.$
Taking the inverse of $c_{i}^{\prime}(\mathcal{X}(\rho)_{i})$ (which exists since $c_{i}$ is strictly convex) yields the desired result. ∎
Proposition 15.
Proof of Proposition 15.
We’ll extend the monotonicity result from Proposition 3 to show that with $\beta$-strongly convex cost functions, the game becomes strongly monotone.
Recall from the proof of Proposition 3 that:
We already established that the first term is non-positive:
Summing over all players:
Therefore:
Proposition 16.
Assume that all the cost functions $c_{i}(x_{i})$ are $\beta$-strongly convex. Then, $\mathcal{X}(\cdot)$ is Lipschitz continuous with respect to the $\ell_{2}$-norm on the interval $[0,1]$, with Lipschitz constant that satisfies $L\leq\frac{\mu\sqrt{n}}{\beta}\frac{X_{UB}}{(X_{LB})^{2-\gamma}}$, where $X_{UB}=\sum_{j=1}^{n}x_{j}^{\max}$ and $X_{LB}=\sum_{i=1}^{n}x_{i}^{\min}$.
Proof of Proposition 16.
Step 1: Characterizing the equilibrium.
Denote $K(\rho)=\frac{\mu(1+\alpha\rho)}{1+\alpha}$. The utility for creator $i$ under enforced full sharing ($\mathbf{s}=\mathbf{1}$) is given by Lemma 9:
The ESE $\mathcal{X}(\rho)$ is characterized by the first-order conditions $\frac{\partial U_{i}}{\partial x_{i}}=0$ for all $i\in[n]$. Let us define $g_{i}(\mathbf{x},\rho)=\frac{\partial U_{i}}{\partial x_{i}}$. Computing this derivative:
$g_{i}(\mathbf{x},\rho)$ $=K(\rho)\frac{\partial}{\partial x_{i}}\left[\left\|\mathbf{x}\right\|^{\gamma-1}x_{i}\right]-c_{i}^{\prime}(x_{i})$
$=K(\rho)\left[(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-2}x_{i}+\left\|\mathbf{x}\right\|^{\gamma-1}\right]-c_{i}^{\prime}(x_{i})$
$=K(\rho)\left\|\mathbf{x}\right\|^{\gamma-2}\left[\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|\right]-c_{i}^{\prime}(x_{i}).$
Let $\mathbf{g}(\mathbf{x},\rho)=(g_{1}(\mathbf{x},\rho),\ldots,g_{n}(\mathbf{x},\rho))^{T}$. By definition, the ESE satisfies $\mathbf{g}(\mathcal{X}(\rho),\rho)=\mathbf{0}$.
Step 2: Continuous differentiability of $\mathcal{X}(\cdot)$.
We verify that $\mathbf{g}(\mathbf{x},\rho)$ is continuously differentiable in both $\mathbf{x}$ and $\rho$:
  • $K(\rho)=\frac{\mu(1+\alpha\rho)}{1+\alpha}$ is clearly $C^{1}$ in $\rho$.
  • From Lemma 12 and the arguments in the proof of Theorem 5, we know that $\mathcal{X}(\rho)_{i}>0$ for all $i\in[n]$, which implies $\left\|\mathcal{X}(\rho)\right\|>0$. Consequently, terms like $\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}$ are well-defined and $C^{1}$ in a neighborhood of $\left\|\mathcal{X}(\rho)\right\|$.
This condition of strong co-monotonicity for a continuously differentiable map $\mathbf{g}$ is equivalent to the symmetric part of its Jacobian, $J_{S}(\mathbf{x},\rho)=\frac{1}{2}(J_{x}(\mathbf{x},\rho)+J_{x}(\mathbf{x},\rho)^{T})$, being negative definite. This, in turn, implies that the Jacobian $J_{x}(\mathcal{X}(\rho),\rho)$ itself is invertible. To see why, assume for contradiction that $J_{x}(\mathcal{X}(\rho),\rho)$ is not invertible. Then, there exists a non-zero vector $\mathbf{u}$ such that $J_{x}(\mathcal{X}(\rho),\rho)\mathbf{u}=\mathbf{0}$. This would lead to:
Step 3: Bounding the derivative norm.
First, we bound $\left\|J_{x}(\mathcal{X}(\rho),\rho)^{-1}\right\|_{2}$. The strong co-monotonicity of $\mathbf{g}(\mathbf{x},\rho)$ implies that for any non-zero vector $\mathbf{v}$:
Lemma 14 ensures that $X_{LB}\leq\left\|\mathcal{X}(\rho)\right\|$. Additionally, Lemma 10 suggests that $\left\|\mathcal{X}(\rho)\right\|\leq X_{UB}$. Using the fact that $\gamma-2<0$, we have $\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}\leq(X_{LB})^{\gamma-2}$. Also, $\gamma\mathcal{X}(\rho)_{i}+\left\|\mathcal{X}(\rho)_{-i}\right\|\leq X_{UB}$ and $\frac{\mu\alpha}{1+\alpha}\leq\mu$. Therefore:
Using the Euclidean norm:
Combining our bounds:
Since $\mathcal{X}(\cdot)$ is continuously differentiable on the compact interval $[0,1]$, its derivative is bounded. By the Mean Value Theorem, for any $\rho_{1},\rho_{2}\in[0,1]$:
$\left\|\mathbf{x}^{\star}(\rho_{1})-\mathbf{x}^{\star}(\rho_{2})\right\|_{2}\leq L\cdot\left|\rho_{1}-\rho_{2}\right|,$
where $L=\sup_{\rho\in[0,1]}\left\|\frac{d\mathbf{x}^{\star}}{d\rho}(\rho)\right\|_{2}\leq\frac{\mu\sqrt{n}}{\beta}\frac{X_{UB}}{(X_{LB})^{2-\gamma}}$.
This establishes that $\mathcal{X}(\cdot)$ is Lipschitz continuous with respect to the $\ell_{2}$-norm on $[0,1]$ with the Lipschitz constant bounded as stated in the proposition, and the proof of Proposition 16 is complete. ∎
@@
Let $Y_{LB}>0$, $Y_{UB}>0$, $x^{\max}>0$, $\alpha>0$, $\mu>0$, and $\gamma\in[0,1]$. Let $c(x)$ be continuously differentiable on $[0,x^{\max}]$ with $\left|c^{\prime}(x)\right|\leq M_{c^{\prime}}$. Define
$U(y)=\max_{x\in[0,x^{\max}]}\left\{\frac{\mu x(x+y)^{\gamma}}{x+(1+\alpha)y}-c(x)\right\},$
and, for $(y,z,\sigma)\in[Y_{LB},Y_{UB}]\times[0,x^{\max}]\times[0,1]$,
$\phi^{0}(y,z,\sigma)=U(y)-\frac{\mu(1+\alpha\sigma)}{1+\alpha}z(z+y)^{\gamma-1}+c(z).$
@@
$\left|\phi^{0}(y_{1},z_{1},\sigma_{1})-\phi^{0}(y_{2},z_{2},\sigma_{2})\right|\leq L_{y,0}^{\star}\left|y_{1}-y_{2}\right|+L_{z,0}^{\star}\left|z_{1}-z_{2}\right|+L_{\sigma,0}^{\star}\left|\sigma_{1}-\sigma_{2}\right|,$
where
$L_{z,0}^{\star}$ $\leq\mu(Y_{LB})^{\gamma-2}(x^{\max}+Y_{UB})+M_{c^{\prime}},$
$L_{y,0}^{\star}$ $\leq\mu x^{\max}(Y_{LB})^{\gamma-2}+\frac{\mu x^{\max}(1+\alpha)}{(Y_{LB})^{3-\gamma}}(x^{\max}+Y_{UB}),$
$L_{\sigma,0}^{\star}$ $\leq\mu x^{\max}(Y_{LB})^{\gamma-1}.$
Proof of Lemma 17.
Define the auxiliary functions
$A(y,z,\sigma)=\frac{\mu(1+\alpha\sigma)}{1+\alpha}z(z+y)^{\gamma-1}-c(z),$
and
$F(x,y)=\frac{\mu x(x+y)^{\gamma}}{x+(1+\alpha)y}-c(x),$
so that $\phi^{0}(y,z,\sigma)=U(y)-A(y,z,\sigma)$, where $U(y)=\max_{x\in[0,x^{\max}]}F(x,y)$.
We bound the sensitivity of $\phi^{0}$ with respect to each variable.
Lipschitz constant in $z$.
$U(y)$ does not depend on $z$. We consider the $z$-derivative of $-A(y,z,\sigma)$:
@@
$=\frac{\mu(1+\alpha\sigma)}{1+\alpha}(z+y)^{\gamma-2}(y+\gamma z)+c^{\prime}(z).$
@@
$\left|\frac{\partial}{\partial z}\left(-A(y,z,\sigma)+c(z)\right)\right|\leq\mu(Y_{LB})^{\gamma-2}(Y_{UB}+x^{\max})+M_{c^{\prime}}\eqqcolon L_{z,0}^{\star}.$
Lipschitz constant in $\sigma$.
@@
$\left|\frac{\partial}{\partial\sigma}\left(-A(y,z,\sigma)\right)\right|\leq\mu x^{\max}(Y_{LB})^{\gamma-1}\eqqcolon L_{\sigma,0}^{\star}.$
Lipschitz constant in $y$.
@@
Next, for $-A(y,z,\sigma)$, the derivative is
$-\frac{\partial A}{\partial y}=-\frac{\mu(1+\alpha\sigma)}{1+\alpha}z(\gamma-1)(z+y)^{\gamma-2}.$
@@
Therefore,
$L_{y,0}^{\star}=\mu x^{\max}(Y_{LB})^{\gamma-2}+\frac{\mu x^{\max}(1+\alpha)}{(Y_{LB})^{3-\gamma}}(x^{\max}+Y_{UB}).$
Conclusion.
For any $(y_{1},z_{1},\sigma_{1}),(y_{2},z_{2},\sigma_{2})$,
$\left|\phi^{0}(y_{1},z_{1},\sigma_{1})-\phi^{0}(y_{2},z_{2},\sigma_{2})\right|\leq L_{y,0}^{\star}\left|y_{1}-y_{2}\right|+L_{z,0}^{\star}\left|z_{1}-z_{2}\right|+L_{\sigma,0}^{\star}\left|\sigma_{1}-\sigma_{2}\right|,$
with the constants as above. This completes the proof of Lemma 17. ∎

### v2

@@ B.2 Proofs of Section 3’s Main Results
Combining these observations, we conclude that no creator $i$ has a profitable deviation from $(\mathcal{X}_{i}(\rho),1)$, making $(\mathcal{X}(\rho),\mathbf{1})$ a PNE. Moreover, uniqueness follows from the observation that any FSE must be an ESE when restricted to the subgame with $\mathbf{s}=\mathbf{1}$. Since Lemma˜3 establishes that the ESE is unique, there can be no other FSE. This completes the proof of Theorem˜5. ∎
See 6
Proof of Proposition˜6.
Fix $w\in(0,1]$. By Theorem˜5, a sufficient condition for $(\mathcal{X}(\rho),\mathbf{1})$ to be the unique FSE is:
$\rho>\max_{i\in[n]}\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|}.$
Since we consider $\rho\in(w,1]$, it suffices to show that for sufficiently large $n$, the RHS is bounded by $w$.
First, we bound the numerator. Recall that $x_{i}^{\max}$ satisfies $a_{i}(x_{i}^{\max})^{\theta}=\mu(x_{i}^{\max})^{\gamma}$. Thus, for all $i$:
$x_{i}^{\max}=\left(\frac{\mu}{a_{i}}\right)^{\frac{1}{\theta-\gamma}}\leq\left(\frac{\mu}{a_{\min}}\right)^{\frac{1}{\theta-\gamma}}\eqqcolon M.$ (10)
Next, we bound the denominator using Lemma˜16, which establishes that $\left\|\mathcal{X}_{-i}(\rho)\right\|=\Theta(n^{\frac{\theta-1}{\theta-\gamma}})$ for any $\rho\in[0,1]$. Specifically, there exists a constant $K>0$ (depending on $\mu,\alpha,\theta,\gamma,a_{\min},a_{\max}$, but not on $n$) such that for all sufficiently large $n$:
$\left\|\mathcal{X}_{-i}(\rho)\right\|\geq K\cdot n^{\frac{\theta-1}{\theta-\gamma}}.$ (11)
Substituting (10) and (11) into the sufficient condition, we require:
$\frac{M}{M+(1+\alpha)Kn^{\frac{\theta-1}{\theta-\gamma}}}\leq w.$
Rearranging to isolate $n$:
$M$ $\leq w\left(M+(1+\alpha)Kn^{\frac{\theta-1}{\theta-\gamma}}\right)$
$M(1-w)$ $\leq w(1+\alpha)Kn^{\frac{\theta-1}{\theta-\gamma}}$
$n^{\frac{\theta-1}{\theta-\gamma}}$ $\geq\frac{M(1-w)}{w(1+\alpha)K}.$
Since $\theta>1$ and $\theta>\gamma$, the exponent $\frac{\theta-1}{\theta-\gamma}$ is positive. Thus, there exists a threshold $N(w)$ such that for all $n>N(w)$, the inequality holds. This ensures the FSE condition is met for all $\rho\in(w,1]$. ∎
See 7
Proof of Proposition˜7.
By symmetry and Lemma˜3, the equilibrium is unique and symmetric, characterized by a single quality level $x(\rho)$–we have $x_{i}=x(\rho)$ for all $i$, which implies $\left\|\mathbf{x}\right\|=nx(\rho)$. Differentiating $U_{i}$ with respect to $x_{i}$ yields:
$\frac{\partial U_{i}}{\partial x_{i}}=\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathbf{x}\right\|^{\gamma-2}\left[(\gamma-1)x_{i}+\left\|\mathbf{x}\right\|\right]-c^{\prime}(x_{i}).$
Setting the derivative to zero and substituting the power cost derivative $c^{\prime}(x)=a\theta x^{\theta-1}$ yields the following FOC (which is enforcing, as Lemma˜15 asserts that all the qualities are positive):
$\frac{\mu(1+\alpha\rho)}{1+\alpha}(nx(\rho))^{\gamma-2}\left[(\gamma-1)x(\rho)+nx(\rho)\right]=a\theta x(\rho)^{\theta-1}.$
Simplifying the term in brackets to $x(\rho)(n+\gamma-1)$ and rearranging to isolate $x(\rho)$, we obtain:
$x(\rho)^{\theta-\gamma}=C\cdot(1+\alpha\rho),$
where $C>0$ is a constant independent of $\rho$. Since $\alpha>0$ and $\theta>1\geq\gamma$ (implying $\theta-\gamma>0$), the right-hand side is strictly increasing in $\rho$. Consequently, $x(\rho)$, and thus total quality $\left\|\mathcal{X}(\rho)\right\|=nx(\rho)$, must be strictly increasing in $\rho$. ∎
@@ Appendix C Omitted Proofs From Section 4
This section is mainly devoted to proving Theorem˜8, along with the other results of Section˜4. The road to the actual proof of Theorem˜8 contains extensive Lipschitz continuity analysis of various auxiliary functions. The structure of this section is as follows:
@@
    We first introduce some notations and prove a lower bound on the quality produced by each creator in any ESE in Section˜C.1.
@@
    Next, in Section˜C.2, we prove several technical Lipschitz continuity results that will be used in subsequent sections.
@@
    In Section˜C.3, we prove Lipschitz continuity results that are directly used in the proof of Algorithm˜1, which build upon the technical results of Section˜C.2.
C.1 ESE Quality Lower Bound and Notations
@@
$x_{i}^{\min}\coloneqq\left(c_{i}^{\prime}\right)^{-1}\left(\frac{\mu(1+\alpha\rho)}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma\right).$
@@
Fix an arbitrary $\rho$. Then, for every creator $i\in[n]$, it holds that $\mathcal{X}_{i}(\rho)\geq x_{i}^{\min}$.
Proof of Lemma˜18.
Fix any creator $i$. Recall that its FOC in the ESE is:
$\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}\left[(\gamma-1)\mathcal{X}_{i}(\rho)+\left\|\mathcal{X}(\rho)\right\|\right]=c_{i}^{\prime}(\mathcal{X}_{i}(\rho)).$
Next, we rearrange the LHS of the FOC and bound it from below:
$c_{i}^{\prime}(\mathcal{X}_{i}(\rho))$ $=\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}\left[(\gamma-1)\mathcal{X}_{i}(\rho)+\left\|\mathcal{X}(\rho)\right\|\right]$
$=\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\left[1+(\gamma-1)\frac{\mathcal{X}_{i}(\rho)}{\left\|\mathcal{X}(\rho)\right\|}\right]$
$\geq\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\gamma$
$\geq\frac{\mu(1+\alpha\rho)}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma,$
Where the last inequality follows from Lemma˜13 and the fact that $\gamma-1\leq 0$. Recall that $c_{i}^{\prime}(\cdot)$ is strictly increasing since $c_{i}(\cdot)$ is strictly convex; thus, we have:
$\mathcal{X}_{i}(\rho)\geq\left(c_{i}^{\prime}\right)^{-1}\left(\frac{\mu(1+\alpha\rho)}{1+\alpha}\left(\sum_{j=1}^{n}x_{j}^{\max}\right)^{\gamma-1}\gamma\right)=x_{i}^{\min}.$
Remark 19.
Throughout this section, we use the following notation:
  • $X_{LB}=\sum_{i=1}^{n}x_{i}^{\min}$ and $X_{UB}=\sum_{j=1}^{n}x_{j}^{\max}$ are lower and upper bounds on the total quality produced in any ESE, respectively.
  • $Y_{LB}=\sum_{i=1}^{n}x_{i}^{\min}-\max_{i\in[n]}x_{i}^{\min}$ and $Y_{UB}=\sum_{j=1}^{n}x_{j}^{\max}-\min_{i\in[n]}x_{i}^{\max}$ are lower and upper bounds on the total quality produced by all creators except any single creator in any ESE, respectively.
  • $x^{\max}=\max_{i\in[n]}x_{i}^{\max}$ is an upper bound on the quality any single creator can produce.
  • $M_{c^{\prime}}=\max_{i\in[n]}\sup_{x\in[0,x^{\max}]}|c_{i}^{\prime}(x)|$ is a uniform bound on the derivatives of all cost functions over the interval $0,x^{\max}]$.
  • $B(q,u,v,w,t)$ is a parametrized bound defined as below for non-negative scalars $q,u,v,w,t$:
    $B(q,u,v,w,t)=q\cdot M_{c^{\prime}}+\frac{\mu\cdot u\cdot n^{v}(1+x^{\max}+Y_{UB})^{w}}{(Y_{LB})^{t}}.$
    Notice that:
    1. 1.
      Increasing both $w$ and $t$ by the same amount increases $B$.
    2. 2.
      $\max\{B(q_{1},u_{1},v_{1},w_{1},t),B(q_{2},u_{2},v_{2},w_{2},t)\}$
      $\leq B(\max\{q_{1},q_{2}\},\max\{u_{1},u_{2}\},\max\{v_{1},v_{2}\},\max\{w_{1},w_{2}\},t)$
  • We will also use the notation $\tilde{B}(q,u,v,w,t)$ to denote a bound of the same form as $B(q,u,v,w,t)$ but with $Y_{LB},Y_{UB}$ replaced by $0.5Y_{LB},2Y_{UB}$, respectively.
C.2 Technical Lipschitz Continuity Results
@@
For $(y,z,\sigma)\in[Y_{LB},Y_{UB}]\times[0,x^{\max}]\times[0,1]$, define the function
$\phi^{0}(y,z,\sigma)=\max_{x\in[0,x^{\max}]}\left\{\frac{\mu x(x+y)^{\gamma}}{x+(1+\alpha)y}-c(x)\right\}-\frac{\mu(1+\alpha\sigma)}{1+\alpha}z(z+y)^{\gamma-1}+c(z).$
@@
$\left|\phi^{0}(y_{1},z_{1},\sigma_{1})-\phi^{0}(y_{2},z_{2},\sigma_{2})\right|$
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2+\alpha,0,2,3-\gamma)\left|y_{1}-y_{2}\right|+B(0,1,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|.$
Proof of Lemma˜20.
Define the following auxiliary functions:
$F(x,y)$ $\coloneqq\frac{\mu x(x+y)^{\gamma}}{x+(1+\alpha)y}-c(x),$
$U(y)$ $\coloneqq\max_{x\in[0,x^{\max}]}F(x,y),$
$A(y,z,\sigma)$ $\coloneqq\frac{\mu(1+\alpha\sigma)}{1+\alpha}z(z+y)^{\gamma-1}-c(z),$
so that $\phi^{0}(y,z,\sigma)=U(y)-A(y,z,\sigma)$. We bound the sensitivity of $\phi^{0}$ with respect to each variable.
Lipschitz constant in $z$
$U(y)$ does not depend on $z$. We consider the $z$-derivative of $A(y,z,\sigma)$:
@@
$=-\frac{\mu(1+\alpha\sigma)}{1+\alpha}(z+y)^{\gamma-2}(y+\gamma z)+c^{\prime}(z).$
@@
$\left|\frac{\partial}{\partial z}\left(-A(y,z,\sigma)\right)\right|\leq\mu(Y_{LB})^{\gamma-2}(Y_{UB}+x^{\max})+M_{c^{\prime}}\leq B(1,1,0,1,2-\gamma)$
Lipschitz constant in $\sigma$
@@
$\left|\frac{\partial}{\partial\sigma}\left(-A(y,z,\sigma)\right)\right|\leq\mu x^{\max}(Y_{LB})^{\gamma-1}\leq B(0,1,0,1,1-\gamma).$
Lipschitz constant in $y$
@@
Next, for $A(y,z,\sigma)$, the derivative is
$\frac{\partial A}{\partial y}=-\frac{\mu(1+\alpha\sigma)}{1+\alpha}z(\gamma-1)(z+y)^{\gamma-2}.$
@@
Therefore, the bound for the Lipschitz constant in $y$ is:
$\mu x^{\max}(Y_{LB})^{\gamma-2}+\frac{\mu x^{\max}(1+\alpha)}{(Y_{LB})^{3-\gamma}}(x^{\max}+Y_{UB})\leq B(0,2+\alpha,0,2,3-\gamma).$
Conclusion
Combining the three Lipschitz constants, we have:
$\left|\phi^{0}(y_{1},z_{1},\sigma_{1})-\phi^{0}(y_{2},z_{2},\sigma_{2})\right|$
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2+\alpha,0,2,3-\gamma)\left|y_{1}-y_{2}\right|+B(0,1,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|.$
This completes the proof of Lemma˜20. ∎